OpenStudy (anonymous):

I got a “No Solution” for the answer. Can someone double check for me to make sure I’m doing the steps correctly? Please Solve for x ln(x)=ln(x-4)+ln(6) Simplifying ln(x) = ln(x + -4) + ln(6) Multiply ln * x lnx = ln(x + -4) + ln(6) Reorder the terms for easier multiplication: lnx = -4ln + lnx + 6ln Reorder the terms: lnx = -4ln + 6ln + lnx Combine like terms: -4ln + 6ln = 2ln lnx = 2ln + lnx Add '-1lnx' to each side of the equation. lnx + -1lnx = 2ln + lnx + -1lnx Combine like terms:2ln + -2ln = 0 -2ln = 0 Divide each side by '-2'.ln = 0 Answer: ln=0

OpenStudy (anonymous):

im back

OpenStudy (anonymous):

u need help

OpenStudy (anonymous):

Just a quick check please. The equation is ln(x)=ln(x-4)+ln6 and I have to solve for x. I came up with a no solution for x.

OpenStudy (anonymous):

ur answer is correct

OpenStudy (anonymous):

You rock!! Thank you

OpenStudy (anonymous):

np!

OpenStudy (anonymous):

\[\ln x = \ln(x-4)+\ln6\] \[\ln x - \ln(x-4)-\ln6\ = 0\] \[\ln \frac{ x }{ 6(x-4) } = 0\] \[e^{\ln \frac{ x }{ 6(x-4) }} = e^{0}\] \[\frac{ x }{ 6(x-4) } = 1\] \[x = \frac{ 24 }{ 5 }\]

OpenStudy (anonymous):

@mathgeek! I already put the answer down as "no solution" is he correct?

OpenStudy (anonymous):

correct! wait let me make sure

OpenStudy (anonymous):

yeah its right!

OpenStudy (anonymous):

He's right or your right O_o

OpenStudy (anonymous):

we both are

OpenStudy (anonymous):

hahaha- so it just depends on how you work the problem right?

OpenStudy (anonymous):

I'm not sure how there can be no solution and x=24/5... eeek

OpenStudy (anonymous):

\[\ln (x)?\] Is that in the equation?

OpenStudy (anonymous):

exactly take it step by step

OpenStudy (anonymous):

ok. I ran it through 2 calculators as a double chekc and 2 said no solution, and 2 came up with 25/5.

OpenStudy (anonymous):

yeah i think your right

OpenStudy (anonymous):

correction 24/5. Guess it just depends on how you work it out... You can either get no solution or 24/5 as an answer- VERY Confusing.

OpenStudy (anonymous):

if its the mathematical ln function then its 24/5 is the right answer!

OpenStudy (anonymous):

very confusing!lol

OpenStudy (anonymous):

Aggghhhhh- this is why I do not do math on purpose.

OpenStudy (raden):

use the property of log, we have to simplify in right side, first ln(x)=ln(x-4)+ln(6) ln(x)=ln{(x-4)6} now, cancel out the ln for both sides we get x = (x-4)6 x = 6x-24 24 = 6x-x 24 = 5x x=24/5

OpenStudy (anonymous):

S@#T guess it's wrong then...

OpenStudy (anonymous):

It is 24/5!!

OpenStudy (anonymous):

yeah good job!

OpenStudy (raden):

yup, in maths dont do more guess! :P

OpenStudy (anonymous):

i just did the problem on paper nd the answer is 24/5

OpenStudy (anonymous):

Emmille- You said that the "no solution" was correct... I'm telling Stanford! lol

OpenStudy (anonymous):

Thanks RadEn

OpenStudy (raden):

you're welcome :)

OpenStudy (anonymous):

ugh Stanford is gonna be pissed

OpenStudy (anonymous):

Yes, yes they are!

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

maybe you'll get the next one- don't fret :)

OpenStudy (anonymous):

lmao

OpenStudy (anonymous):

need anymore help before i log off